# The Seifert-van Kampen Theorem via Computational Paths: A Formalized Approach to Computing Fundamental Groups

### v1

@@ 4.1. Free Product Words
{defi}
[Free Product Word] A word in the free product $G_{1}\ast G_{2}$ is an alternating sequence of elements from $G_{1}$ and $G_{2}$:
inductive FreeProductWord (G1 : Type u) (G2 : Type u) : Type u
| nil : FreeProductWord G1 G2
| consLeft (x : G1) (rest : FreeProductWord G1 G2) : FreeProductWord G1 G2
| consRight (y : G2) (rest : FreeProductWord G1 G2) : FreeProductWord G1 G2
@@
def concat : FreeProductWord G1 G2 -> FreeProductWord G1 G2
-> FreeProductWord G1 G2
| nil, w2 => w2
| consLeft x rest, w2 => consLeft x (concat rest w2)
| consRight y rest, w2 => consRight y (concat rest w2)
Lemma 3 (Word Concatenation Properties).
@@ 4.2. Amalgamated Free Product
{defi}
[Amalgamation Relation] The amalgamation relation identifies elements from $H$ mapped into $G_{1}$ with their counterparts mapped into $G_{2}$:
inductive AmalgRelation {G1 G2 H : Type u}
(i1 : H -> G1) (i2 : H -> G2) :
FreeProductWord G1 G2 -> FreeProductWord G1 G2 -> Prop
| amalgLeftToRight (h : H) (pre suf : FreeProductWord G1 G2) :
AmalgRelation i1 i2
(concat pre (concat (singleLeft (i1 h)) suf))
(concat pre (concat (singleRight (i2 h)) suf))
| amalgRightToLeft (h : H) (pre suf : FreeProductWord G1 G2) :
AmalgRelation i1 i2
(concat pre (concat (singleRight (i2 h)) suf))
(concat pre (concat (singleLeft (i1 h)) suf))
{defi}
[Amalgamated Free Product] The amalgamated free product is the quotient of free product words by the equivalence closure of the amalgamation relation:
def AmalgamatedFreeProduct (G1 G2 H : Type u)
(i1 : H -> G1) (i2 : H -> G2) : Type u :=
Quot (AmalgEquiv i1 i2)
@@
Remark 4 (Structure of AmalgEquiv).
@@
    The standard word-reduction equivalences for the free product $G_{1}\ast G_{2}$: combining adjacent letters from the same factor (e.g., $\mathrm{consLeft}(x,\mathrm{consLeft}(y,w))\sim\mathrm{consLeft}(x\cdot y,w)$), and removing identity elements.
@@
In the Lean formalization, we build the free product quotient first using standard group-theoretic reductions, then extend with the amalgamation relations. The result is the pushout in the category of groups.
Lemma 5 (Amalgamation Property).
@@
Proof 4.1.
By definition, the singleton words $\mathrm{singleLeft}(i_{1}(h))$ and $\mathrm{singleRight}(i_{2}(h))$ are related by $\mathrm{AmalgRelation}$ (with empty prefix and suffix). The quotient $\mathrm{Quot.sound}$ then yields the equality.
@@ 5. The Seifert-van Kampen Theorem
With pushouts (Section 3) and amalgamated free products (Section 4) in hand, we can now state and prove the main theorem. The key is to establish a bijective correspondence between loops in the pushout and words in the amalgamated free product. The encode-decode method provides a systematic approach: we construct maps in both directions and prove they are mutually inverse.
@@ 5.1. Statement
Theorem 6 (Seifert-van Kampen).
@@ 5.2.1. The Decode Map
{defi}
[Pushout Decode] The decode map converts a word to a loop in the pushout:
noncomputable def pushoutDecode (c0 : C) :
FreeProductWord (pi_1(A, f c0)) (pi_1(B, g c0))
-> pi_1(Pushout A B C f g, Pushout.inl (f c0))
| .nil => Quot.mk _ (Path.refl _)
| .consLeft alpha rest =>
-- alpha : pi_1(A), map to pi_1(Pushout) via inl
let alphaInPushout := Quot.lift
(fun p => Quot.mk RwEq (Pushout.inlPath p))
(fun _ _ hp => Quot.sound (rweq_congrArg_of_rweq Pushout.inl hp))
alpha
piOneMul alphaInPushout (pushoutDecode c0 rest)
| .consRight beta rest =>
-- beta : pi_1(B), needs transport via glue
let betaInPushout := Quot.lift
(fun q => Quot.mk RwEq (Path.trans
(Pushout.glue c0)
(Path.trans (Pushout.inrPath q)
(Path.symm (Pushout.glue c0)))))
(fun _ _ hq => Quot.sound (...))
piOneMul betaInPushout (pushoutDecode c0 rest)
@@
$\beta\mapsto\mathrm{glue}(c_{0})\cdot\mathrm{inrPath}(\beta)\cdot\mathrm{glue}(c_{0})^{-1}$
@@ 5.2.2. Decode Respects Amalgamation
Lemma 7 (Decode Respects Amalgamation).
@@
$\mathrm{inlPath}(f_{*}(p))\sim\mathrm{glue}(c_{0})\cdot\mathrm{inrPath}(g_{*}(p))\cdot\mathrm{glue}(c_{0})^{-1}$
This follows from glue naturality (Lemma 2) applied to the loop $p$:
theorem glue_natural_loop_rweq (c0 : C) (p : LoopSpace C c0) :
RwEq (inlPath (Path.congrArg f p))
(Path.trans (glue c0)
(Path.trans (inrPath (Path.congrArg g p))
(Path.symm (glue c0)))) :=
glue_natural_rweq p
@@ 5.2.3. The Encode Map
The encode map uses the “code family” approach, a technique central to HoTT calculations of fundamental groups [10, 16]. The idea is to define a type family $\mathrm{Code}:\mathrm{Pushout}\to\mathrm{Type}$ that, at each point, represents the words that “should decode to a path ending at that point.” This family is defined via the pushout recursion principle:
  • At $\mathrm{inl}(a)$: $\mathrm{Code}(\mathrm{inl}(a))$ represents words that decode to paths ending at $\mathrm{inl}(a)$
  • At $\mathrm{inr}(b)$: $\mathrm{Code}(\mathrm{inr}(b))$ represents words that decode to paths ending at $\mathrm{inr}(b)$
  • Transport along $\mathrm{glue}(c)$: adjusts the word by prepending or removing letters corresponding to the glue path, ensuring coherence across the pushout
The encode function is then defined as transport of the empty word along a path:
@@
For a loop $p:\mathrm{Path}(\mathrm{inl}(f(c_{0})),\mathrm{inl}(f(c_{0})))$, this yields a word representing the “symbolic content” of $p$.
@@
Fully constructing the code family requires dependent elimination over the pushout HIT, which in Lean 4 (without native HIT support or cubical features) would require extensive additional infrastructure. Since our focus is on the computational paths framework rather than the HIT metatheory, we axiomatize encode’s existence and key properties:
axiom pushoutEncodeAxiom (A B C : Type u) (f : C -> A) (g : C -> B) (c0 : C) :
LoopSpace (Pushout A B C f g) (Pushout.inl (f c0)) ->
FreeProductWord (pi_1(A, f c0)) (pi_1(B, g c0))
axiom pushoutEncodeAxiom_respects_rweq (A B C : Type u)
(f : C -> A) (g : C -> B) (c0 : C)
{p q : LoopSpace (Pushout A B C f g) (Pushout.inl (f c0))}
(h : RwEq p q) :
pushoutEncodeAxiom A B C f g c0 p = pushoutEncodeAxiom A B C f g c0 q
The second axiom ensures that encode respects rewrite equality, which is essential for lifting encode to the fundamental group quotient. In a cubical type theory setting, these axioms would be provable; here, they capture exactly the interface needed for the SVK proof.
@@ 5.2.4. Round-Trip Properties
Lemma 8 (Decode-Encode Round-Trip).
@@
Lemma 9 (Encode-Decode Round-Trip).
@@ 5.3. The Equivalence
noncomputable def seifertVanKampenEquiv :
SimpleEquiv
(pi_1(Pushout A B C f g, Pushout.inl (f c0)))
(AmalgamatedFreeProduct
(pi_1(A, f c0))
(pi_1(B, g c0))
(pi_1(C, c0))
(piOneFmap c0)
(piOneGmap c0)) where
toFun := pushoutEncodeAmalg c0
invFun := pushoutDecodeAmalg c0
left_inv := by
intro alpha
induction alpha using Quot.ind with
| _ p => exact pushoutDecodeEncodeAxiom A B C f g c0 p
right_inv := by
intro x
induction x using Quot.ind with
| _ w =>
apply Quot.sound
exact pushoutEncodeDecodeAxiom A B C f g c0 w
@@ 6.1. The Wedge Sum: Free Product of Fundamental Groups
Theorem 10 (Fundamental Group of Wedge Sum).
@@
Proof 6.1.
The wedge $A\vee B=\mathrm{Pushout}(A,B,\mathrm{Unit},\lambda\_.a_{0},\lambda\_.b_{0})$. By SVK:
$\pi_{1}(A\vee B)\simeq\pi_{1}(A)\ast_{\pi_{1}(\mathrm{Unit})}\pi_{1}(B)$
Since $\pi_{1}(\mathrm{Unit})\cong 1$, the amalgamation is trivial and we get the free product.
@@
noncomputable def wedgeFundamentalGroupEquiv :
SimpleEquiv
(pi_1(Wedge A B a0 b0, Wedge.basepoint))
(FreeProductWord (pi_1(A, a0)) (pi_1(B, b0))) where
toFun := wedgeEncodeQuot a0 b0
invFun := wedgeFreeProductDecode a0 b0
left_inv := ...
right_inv := ...
@@ 6.2. The Figure-Eight Space
Theorem 11 (Fundamental Group of Figure-Eight).
@@
using the circle theorem $\pi_{1}(S^{1})\simeq\mathbb{Z}$.
@@
def FigureEight : Type u := Wedge Circle Circle circleBase circleBase
noncomputable def figureEightPiOneEquivPiOneProduct :
SimpleEquiv FigureEightPiOne (FreeProductWord circlePiOne circlePiOne) :=
wedgeFundamentalGroupEquiv circleBase circleBase
noncomputable def figureEightPiOneEquiv :
SimpleEquiv FigureEightPiOne (FreeProductWord Int Int) :=
SimpleEquiv.comp figureEightPiOneEquivPiOneProduct piOneProductEquivIntProduct
@@ 6.2.1. Non-Abelianness
def wordAB : FreeProductWord Int Int := .consLeft 1 (.consRight 1 .nil)
def wordBA : FreeProductWord Int Int := .consRight 1 (.consLeft 1 .nil)
theorem wordAB_ne_wordBA : wordAB != wordBA := by
intro h
cases h
This shows that $\mathrm{loop}_{A}\cdot\mathrm{loop}_{B}\neq\mathrm{loop}_{B}\cdot\mathrm{loop}_{A}$. Topologically, $\mathrm{wordAB}$ corresponds to the loop that first traverses the left circle and then the right circle, while $\mathrm{wordBA}$ traverses them in the opposite order. In the figure-eight, these loops cannot be continuously deformed into each other—they are genuinely distinct homotopy classes. This non-commutativity is the hallmark of the free product structure and explains why the figure-eight has a more complex fundamental group than a single circle or a torus.

### v2

@@ 4.1. Free Product Words
Definition 4.1 (Free Product Word).
A word in the free product $G_{1}\ast G_{2}$ is an alternating sequence of elements from $G_{1}$ and $G_{2}$:
1inductive FreeProductWord (G1 : Type u) (G2 : Type u) : Type u
2 | nil : FreeProductWord G1 G2
3 | consLeft (x : G1) (rest : FreeProductWord G1 G2) : FreeProductWord G1 G2
4 | consRight (y : G2) (rest : FreeProductWord G1 G2) : FreeProductWord G1 G2
Example 4.2 (Figure-Eight Words).
For $\pi_{1}(S^{1}\vee S^{1})\simeq\mathbb{Z}\ast\mathbb{Z}$, words represent sequences of loops around the two circles:
  • $\mathrm{consLeft}(1,\mathrm{nil})$ — one loop around the left circle
  • $\mathrm{consRight}(2,\mathrm{nil})$ — two loops around the right circle
  • $\mathrm{consLeft}(1,\mathrm{consRight}(-1,\mathrm{nil}))$ — left loop, then reverse right loop
@@
1def concat : FreeProductWord G1 G2 -> FreeProductWord G1 G2
2 -> FreeProductWord G1 G2
3 | nil, w2 => w2
4 | consLeft x rest, w2 => consLeft x (concat rest w2)
5 | consRight y rest, w2 => consRight y (concat rest w2)
Lemma 4.3 (Word Concatenation Properties).
@@
Free vs. Direct Product.
Unlike the torus $S^{1}\times S^{1}$ whose loops can “slide past each other” (yielding the abelian group $\mathbb{Z}\times\mathbb{Z}$), loops in $S^{1}\vee S^{1}$ must pass through the common basepoint, preventing commutation. The word $\mathrm{loop}_{A}\cdot\mathrm{loop}_{B}$ represents first traversing the left circle, then the right; reversing the order gives a genuinely distinct homotopy class $\mathrm{loop}_{B}\cdot\mathrm{loop}_{A}$.
@@ 4.2. Amalgamated Free Product
Definition 4.4 (Amalgamation Relation).
The amalgamation relation identifies elements from $H$ mapped into $G_{1}$ with their counterparts mapped into $G_{2}$:
1inductive AmalgRelation {G1 G2 H : Type u}
2 (i1 : H -> G1) (i2 : H -> G2) :
3 FreeProductWord G1 G2 -> FreeProductWord G1 G2 -> Prop
4 | amalgLeftToRight (h : H) (pre suf : FreeProductWord G1 G2) :
5 AmalgRelation i1 i2
6 (concat pre (concat (singleLeft (i1 h)) suf))
7 (concat pre (concat (singleRight (i2 h)) suf))
8 | amalgRightToLeft (h : H) (pre suf : FreeProductWord G1 G2) :
9 AmalgRelation i1 i2
10 (concat pre (concat (singleRight (i2 h)) suf))
11 (concat pre (concat (singleLeft (i1 h)) suf))
Definition 4.5 (Amalgamated Free Product).
The amalgamated free product is the quotient of free product words by the equivalence closure of the amalgamation relation:
1def AmalgamatedFreeProduct (G1 G2 H : Type u)
2 (i1 : H -> G1) (i2 : H -> G2) : Type u :=
3 Quot (AmalgEquiv i1 i2)
@@
Remark 4.6 (Structure of AmalgEquiv).
@@
    Standard word-reduction equivalences for the free product $G_{1}\ast G_{2}$: combining adjacent letters from the same factor, and removing identity elements.
@@
In the Lean formalization, we build the free product quotient first, then extend with amalgamation. The result is the pushout in the category of groups.
Lemma 4.7 (Amalgamation Property).
@@
Proof.
By definition, the singleton words $\mathrm{singleLeft}(i_{1}(h))$ and $\mathrm{singleRight}(i_{2}(h))$ are related by $\mathrm{AmalgRelation}$ (with empty prefix and suffix). The quotient $\mathrm{Quot.sound}$ then yields the equality. ∎
@@ 5. The Seifert-van Kampen Theorem
With pushouts (Section 3) and amalgamated free products (Section 4) in hand, we can now state and prove the main theorem.
@@ 5.1. Statement
Theorem 5.1 (Seifert-van Kampen).
@@
Proof sketch.
We construct:
  1. (1)
    $\mathrm{decode}$: Maps words to loops by interpreting left letters as loops in $\mathrm{inl}(A)$, and right letters as loops in $\mathrm{inr}(B)$ conjugated by $\mathrm{glue}(c_{0})$.
  2. (2)
    $\mathrm{encode}$: Defined via transport of a code family along paths (axiomatized).
The key lemmas are:
  • Decode respects amalgamation (Lemma 5.3): elements from $\pi_{1}(C)$ decode identically whether placed in the left or right factor.
  • Round-trip properties: $\mathrm{decode}\circ\mathrm{encode}=\mathrm{id}$ and $\mathrm{encode}\circ\mathrm{decode}\sim_{\mathrm{amalg}}\mathrm{id}$.
The first round-trip follows by path induction on pushout paths; the second by word induction. See §5.2.35.2.4 for details. ∎
@@ 5.2.1. The Decode Map
Definition 5.2 (Pushout Decode).
The decode map converts a word to a loop in the pushout:
1noncomputable def pushoutDecode (c0 : C) :
2 FreeProductWord (pi_1(A, f c0)) (pi_1(B, g c0))
3 -> pi_1(Pushout A B C f g, Pushout.inl (f c0))
4 | .nil => Quot.mk _ (Path.refl _)
5 | .consLeft alpha rest =>
6 let alphaInPushout := Quot.lift
7 (fun p => Quot.mk RwEq (Pushout.inlPath p))
8 (fun _ _ hp => Quot.sound (rweq_congrArg_of_rweq Pushout.inl hp))
9 alpha
10 piOneMul alphaInPushout (pushoutDecode c0 rest)
11 | .consRight beta rest =>
12 let betaInPushout := Quot.lift
13 (fun q => Quot.mk RwEq (Path.trans
14 (Pushout.glue c0)
15 (Path.trans (Pushout.inrPath q)
16 (Path.symm (Pushout.glue c0)))))
17 (fun _ _ hq => Quot.sound (...))
18 beta
19 piOneMul betaInPushout (pushoutDecode c0 rest)
@@ 2 FreeProductWord (pi_1(A, f c0)) (pi_1(B, g c0))
$\beta\mapsto\mathrm{glue}(c_{0})\cdot\mathrm{inr}_{*}(\beta)\cdot\mathrm{glue}(c_{0})^{-1}$
$\mathrm{inl}(f(c_{0}))$$\mathrm{inr}(g(c_{0}))$$\mathrm{inl}(f(c_{0}))$$\mathrm{inr}(g(c_{0}))$$\scriptstyle{\mathrm{glue}(c_{0})}$$\scriptstyle{\beta_{\mathrm{decoded}}}$$\scriptstyle{\mathrm{inr}_{*}(\beta)}$$\scriptstyle{\mathrm{glue}(c_{0})^{-1}}$
@@ 5.2.2. Decode Respects Amalgamation
Lemma 5.3 (Decode Respects Amalgamation).
@@
$\mathrm{inl}_{*}(f_{*}(p))\sim\mathrm{glue}(c_{0})\cdot\mathrm{inr}_{*}(g_{*}(p))\cdot\mathrm{glue}(c_{0})^{-1}$
This is exactly Corollary 3.4 applied to the loop $p$ (where $c_{1}=c_{2}=c_{0}$):
1theorem glue_natural_loop_rweq (c0 : C) (p : LoopSpace C c0) :
2 RwEq (inlPath (Path.congrArg f p))
3 (Path.trans (glue c0)
4 (Path.trans (inrPath (Path.congrArg g p))
5 (Path.symm (glue c0)))) :=
6 glue_natural_rweq p
@@ 5.2.3. The Encode Map
The encode map uses the “code family” approach, a technique central to HoTT calculations of fundamental groups [10, 17]. The idea is to define a type family $\mathrm{Code}:\mathrm{Pushout}\to\mathrm{Type}$ that, at each point, represents the words that “should decode to a path ending at that point.”
  • At $\mathrm{inl}(a)$: $\mathrm{Code}(\mathrm{inl}(a))$ represents words decoding to paths ending at $\mathrm{inl}(a)$
  • At $\mathrm{inr}(b)$: $\mathrm{Code}(\mathrm{inr}(b))$ represents words decoding to paths ending at $\mathrm{inr}(b)$
  • Transport along $\mathrm{glue}(c)$: adjusts words by prepending/removing letters for the glue path
The encode function is then:
@@
Fully constructing the code family requires dependent elimination over the pushout HIT, which in Lean 4 (without native HIT support or cubical features) would require extensive additional infrastructure. Since our focus is on the computational paths framework, we axiomatize encode’s existence and key properties:
1axiom pushoutEncodeAxiom (A B C : Type u) (f : C -> A) (g : C -> B) (c0 : C) :
2 LoopSpace (Pushout A B C f g) (Pushout.inl (f c0)) ->
3 FreeProductWord (pi_1(A, f c0)) (pi_1(B, g c0))
4
5axiom pushoutEncodeAxiom_respects_rweq (A B C : Type u)
6 (f : C -> A) (g : C -> B) (c0 : C)
7 {p q : LoopSpace (Pushout A B C f g) (Pushout.inl (f c0))}
8 (h : RwEq p q) :
9 pushoutEncodeAxiom A B C f g c0 p = pushoutEncodeAxiom A B C f g c0 q
The second axiom ensures encode respects $\sim$, essential for lifting to the $\pi_{1}$ quotient. In cubical type theory, these would be provable; here, they capture the interface needed for SVK.
@@ 5.2.4. Round-Trip Properties
Lemma 5.4 (Decode-Encode Round-Trip).
@@
Lemma 5.5 (Encode-Decode Round-Trip).
@@ 5.3. The Equivalence
1noncomputable def seifertVanKampenEquiv :
2 SimpleEquiv
3 (pi_1(Pushout A B C f g, Pushout.inl (f c0)))
4 (AmalgamatedFreeProduct
5 (pi_1(A, f c0))
6 (pi_1(B, g c0))
7 (pi_1(C, c0))
8 (piOneFmap c0)
9 (piOneGmap c0)) where
10 toFun := pushoutEncodeAmalg c0
11 invFun := pushoutDecodeAmalg c0
12 left_inv := by
13 intro alpha
14 induction alpha using Quot.ind with
15 | _ p => exact pushoutDecodeEncodeAxiom A B C f g c0 p
16 right_inv := by
17 intro x
18 induction x using Quot.ind with
19 | _ w =>
20 apply Quot.sound
21 exact pushoutEncodeDecodeAxiom A B C f g c0 w
@@ 6.1. The Wedge Sum: Free Product of Fundamental Groups
Theorem 6.1 (Fundamental Group of Wedge Sum).
@@
Proof.
The wedge $A\vee B=\mathrm{Pushout}(A,B,\mathbf{1},\lambda\_.a_{0},\lambda\_.b_{0})$. By SVK:
$\pi_{1}(A\vee B)\simeq\pi_{1}(A)\ast_{\pi_{1}(\mathbf{1})}\pi_{1}(B)$
Since $\pi_{1}(\mathbf{1})\cong 1$ (the trivial group), the amalgamation imposes no relations, yielding the free product. ∎
@@
1noncomputable def wedgeFundamentalGroupEquiv :
2 SimpleEquiv
3 (pi_1(Wedge A B a0 b0, Wedge.basepoint))
4 (FreeProductWord (pi_1(A, a0)) (pi_1(B, b0))) where
5 toFun := wedgeEncodeQuot a0 b0
6 invFun := wedgeFreeProductDecode a0 b0
7 left_inv := ...
8 right_inv := ...
@@ 6.2. The Figure-Eight Space
Theorem 6.2 (Fundamental Group of Figure-Eight).
@@
using the fundamental theorem $\pi_{1}(S^{1})\simeq\mathbb{Z}$ [10]. ∎
@@
1def FigureEight : Type u := Wedge Circle Circle circleBase circleBase
2
3noncomputable def figureEightPiOneEquivPiOneProduct :
4 SimpleEquiv FigureEightPiOne (FreeProductWord circlePiOne circlePiOne) :=
5 wedgeFundamentalGroupEquiv circleBase circleBase
6
7noncomputable def figureEightPiOneEquiv :
8 SimpleEquiv FigureEightPiOne (FreeProductWord Int Int) :=
9 SimpleEquiv.comp figureEightPiOneEquivPiOneProduct piOneProductEquivIntProduct
@@ 6.2.1. Non-Abelianness
1def wordAB : FreeProductWord Int Int := .consLeft 1 (.consRight 1 .nil)
2def wordBA : FreeProductWord Int Int := .consRight 1 (.consLeft 1 .nil)
3
4theorem wordAB_ne_wordBA : wordAB != wordBA := by
5 intro h
6 cases h -- consLeft and consRight are distinct constructors
This shows that $\mathrm{loop}_{A}\cdot\mathrm{loop}_{B}\neq\mathrm{loop}_{B}\cdot\mathrm{loop}_{A}$. The proof is by case analysis: FreeProductWord is an inductive type where consLeft and consRight are distinct constructors, so no sequence of reductions can make them equal.
Topologically, $\mathrm{wordAB}$ corresponds to first traversing the left circle, then the right; $\mathrm{wordBA}$ traverses them in the opposite order. In the figure-eight, these loops cannot be continuously deformed into each other—they are genuinely distinct homotopy classes.